40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 19 18:10:10 2022, Paco, Update, BHD, Part V of BHR upgrade - POX11 path, LO path, and ITMX Oplev DJI_0167.JPGDJI_0169.JPGDJI_0168.JPG
    Reply  Wed Apr 20 08:20:22 2022, Paco, Update, BHD, Part V of BHR upgrade - PR2 weirdness 
       Reply  Wed Apr 20 18:50:03 2022, Koji, Update, BHD, Part V of BHR upgrade - PR2 weirdness Screenshot_2022-04-20_18-46-11.png
    Reply  Thu Apr 21 16:49:01 2022, Paco, Update, BHD, Part V of BHR upgrade - BS Oplev + LO1 offset 
       Reply  Thu Apr 21 18:42:47 2022, Tega, Update, BHD, Part V of BHR upgrade - Align LO and AS beams to BHD BS 
          Reply  Thu Apr 21 20:33:31 2022, Koji, Update, BHD, AS1 UR OSEM problem localized in the chamber Screenshot_from_2022-04-21_20-13-13.png
             Reply  Fri Apr 22 12:03:09 2022, Tega, Update, BHD, AS1 UR OSEM problem localized in the chamber 
                Reply  Fri Apr 22 12:15:08 2022, Anchal, Update, BHD, Cable post installation 
                   Reply  Mon Apr 25 16:57:57 2022, Anchal, Update, BHD, Cable post installation 
Message ID: 16810     Entry time: Mon Apr 25 16:57:57 2022     In reply to: 16805
Author: Anchal 
Type: Update 
Category: BHD 
Subject: Cable post installation 

[Anchal, Tega, JC]

We installed cable posts in ITMY, BS, and ITMX chambers for all the new suspensions. Now, there is no point where the OSEM connections are hanging freely.

In BS chamber, we installed one post for LO2 near the north edge of the table and another post for PR3 on the Western edge with the blue cable running around the table on the floor.

In ITMY chamber, we installed the cable post in between AS1 and AS4 with the blue cables running around the table on the floor. This is to ensure the useful part of the table remains empty for future and none of the OSEM cables are taught in air.

 

ELOG V3.1.3-